OpenStudy (anikate):

Find the 7th term of the sequence given r=8 and a1= -2

OpenStudy (anonymous):

Wait, geometric sequence: an= a1 r^(n-1)

OpenStudy (anikate):

ok

OpenStudy (anonymous):

Where r represents the common ratio between every number. So, if you think about it, if you multiply 8 to -2 you get -16 8*-2=-16 Notice that 6 is a2. a2=-16 Then, -16*(8)=-128 a3=-128 and you could do this until you get to a7. But, I'm sure you notice that method is tedious. So, applying the nth term of a geometric sequence: an=a1r^(n-1) a7= -2(8)^(7-1) a7=-2(8)^6 a7=-524288

OpenStudy (anonymous):

I hope that helps. It's also helpful to know that a geometric sequence is defined as: A sequence whose consecutive terms have a common ratio. \[\frac{ a2 }{ a1 }=r, \frac{ a3 }{a2 }=r, \frac{ a4 }{ a3 }=r... \] r is the common ratio

OpenStudy (anikate):

oh i got it. thx bro